OpenStudy (anonymous):

Simplify and write in equivalent form with positive exponents. (4x^-2)^-2 So I have my answer, but want to make sure i am on the right track.

OpenStudy (lgbasallote):

what's your answer?

OpenStudy (anonymous):

my answer is -16x^4

OpenStudy (lgbasallote):

close...very..

OpenStudy (lgbasallote):

\[\Large (4x^{-2})^{-2} \implies \frac{1}{(4x^{-2})^2} \implies \frac{1}{16x^{-4}} \implies \frac{x^4}{16}\]

OpenStudy (lgbasallote):

do you get it?

OpenStudy (anonymous):

what I am having trouble understanding is why we add a one in the numberator. Are we dividing?

OpenStudy (lgbasallote):

well i cant just leave the numerator blank can i.. \[\Large \frac{}{(4x^{-2})^2}\] that looks awkward

OpenStudy (lgbasallote):

that's why there's a 1 in the numerator

OpenStudy (lgbasallote):

because 1 is like the identity number

OpenStudy (anonymous):

ok I guess that wouldn't make sense lol.

OpenStudy (lgbasallote):

anything you multiply by 1 is the number itself

OpenStudy (lgbasallote):

maybe this can help you \[\huge a^{-m} = \frac{1}{a^m}\] that means if a number has a negative exponent then you put it in the denominator

OpenStudy (anonymous):

ok, so I do understand that anything multiplied by one is the same number.

OpenStudy (lgbasallote):

so what dont you understand?

OpenStudy (anonymous):

no, actually when you gave me that example, it does make sense. (if I can just remember to do this myself with similar problems.) thank you

OpenStudy (lgbasallote):

ahh. welcome ^_^